the probability that a person has a checking account is 0.74 a saving account is 0.31 and both account is 0.22. Find probability that a randomly selected person has a checking or saving accounts


1
Expert's answer
2021-09-21T06:15:27-0400

"P(C\cup S)=P(C)+P(S)-P(C\cap S)=0.74+0.31-0.22=0.83."